OpenStudy (anonymous):

Solve the following equations algebraically to determine exact solutions over the interval [0.2pi] a) sin 2x +sin 4x = 0 b) cos 2x +sin x =0

OpenStudy (anonymous):

Part (b) first

OpenStudy (anonymous):

cos (2x) = 1 - 2 sin^2 (x)..thats one of the formulas on your formula sheet

OpenStudy (anonymous):

so, 1 - 2 sin^2 (x) + sin x = 0

OpenStudy (anonymous):

Or, 2 sin^2 (x) - sin x - 1 = 0

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now factor, (2sinx + 1)(sinx - 1) = 0

OpenStudy (anonymous):

Set each factor equal to zero, the first factor (2 sin x + 1) = 0, so sin x = - 1/2, that occurs at 210 and 330 degrees, or - 7pi/6 and 11pi/6 radians.

OpenStudy (anonymous):

The second factor, (sin x - 1) = 0 means sin x = 1, and that occurs only at 90 degrees, or pi/2 radians.

OpenStudy (anonymous):

The second factor, (sin x - 1) = 0 means sin x = 1, and that occurs only at 90 degrees, or pi/2 radians.

OpenStudy (anonymous):

Final answers to part(b) equation is pi/2, 7pi/6, and 11pi/6. I think I typed -7pi/6 earlier, thats not right, I meant 7pi/6

OpenStudy (anonymous):

We are now done with part (b). What are your questions in this part (b)?

OpenStudy (anonymous):

When you went from 1-2sin^2x + sinx to 2sin^2x - sinx -1 How did you do that?

OpenStudy (anonymous):

You can change the signs of everything...why do I want to change the signs? Because I want to factor something that starts with 2 sin^2 (x) and NOT - 2sin^2(x).

OpenStudy (anonymous):

So you can change the signs as long as you change every single one.

OpenStudy (anonymous):

Yes, take a look for a moment. We had 1 - 2 sin^2(x) + sin x = 0 So, go ahead and bring everything from the left side to the right side, what do you get? You get 2 sin^2(x) - sin x - 1 = 0 Agree?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So you can bring slowly everything from the left side to the right side, OR you can change all the signs, and leave the left side as it is (with the signs changed).

OpenStudy (anonymous):

I just changed all the signs, and left the left side where it was, with just the signs changed.

OpenStudy (anonymous):

Alright, I see that.

OpenStudy (anonymous):

Do you also see, how I got 7pi/6 and 11pi/6 and pi/2 as the solutions? Thats critical.

OpenStudy (anonymous):

We wanted, as stated in the directions, all angles x in the closed interval [0,2pi].

OpenStudy (anonymous):

Yes. By using the unit circle, and plugging in the "coordinates" of sin (or cos)

OpenStudy (anonymous):

ok. Part (b) was not simple. Part (a) is much trickier, so I left that for last. So I want to take a 2-minute break and will then return to show you part (a). I just need to get a drink.

OpenStudy (anonymous):

Ok :)

OpenStudy (anonymous):

Will be back in approx 2 min

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

I'm back. Ready when you are.

OpenStudy (anonymous):

The equastion in part (a) is sin(2x) + sin (4x) = 0.

OpenStudy (anonymous):

Ready?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

OK. Now, we have a formula for the sin(2x) = 2 sin (x) cos(x). We do NOT have a formula for sin (4x), although we can develop that formula, but I do not think your teacher would want you to take that approach. I will give you a solution which I feel your teacher would want you to present.

OpenStudy (anonymous):

In our equation, let u= 2x. Si if u = 2x, then 2u = 4x. Agree?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So our equation becomes sin (u) + sin (2u) = 0 Agree? I will wait for you.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I got it.

OpenStudy (anonymous):

sin (2u) = 2 sin(u) cos (u)

OpenStudy (anonymous):

so our equation is now sin (u) + 2 sin(u)cos(u) = 0

OpenStudy (anonymous):

Factor out sin (u), we get sin (u) [1 + 2 cos (u)] = 0

OpenStudy (anonymous):

set each factor equal to zero, so sin (u) = 0 AND 2 cos(u) = -1 Or, sin (u) = 0 AND cos (u) = -1/2

OpenStudy (anonymous):

Agree so far?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Now, let's take the first factor and set equal to zero and solve. sin (u) = 0, then u = 0, pi, and 2pi BUT we want to solve for x, not u, so we substitute u = 2x therefore, 2x = 0, 2x = pi, and 2x = 2pi hence, x = 0, pi/2,pi, 3pi/2, 2pi

OpenStudy (anonymous):

We take the second factor, cos (u) = -1/2, and u = 2pi/3, 4pi/3, etc.but u = 2x, so x = pi/3, 2pi/3, pi,4pi/3, and 5pi/3

OpenStudy (anonymous):

Done. So there are many solutions. Now, look everything over and tell me what you dont follow, if anything.

OpenStudy (anonymous):

There?

OpenStudy (anonymous):

Tell me if I made any error, as I do all of this in my head.

OpenStudy (anonymous):

How did you get 3pi/2 and 2pi for the first factor?

OpenStudy (anonymous):

We want all solutions in the closed interval {0,2pi]. Look at the first 3 solutions there, o, pi/2, pi...so every pi/2 theres another solution..so i added pi/2 to pi to get 3pi/2, then i added another pi/2 to 3pi/2 to get 2pi

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Go through each step and you'll see that each step makes sense. It is a tough problem. But the concept of letting u = 2x and 2u = 4x is critical here.

OpenStudy (anonymous):

Feel free to get back to me if any step is not crystal clear.

OpenStudy (anonymous):

Thank you! I'll definitely do that (:

OpenStudy (anonymous):

No problem. Your welcome.